OpenStudy (anonymous):

write an equation in point slope form passes through the points (3,5) and (2,8)? write an equation in point slope form passes through the points (3,5) and (2,8)? @Mathematics

OpenStudy (lalaly):

do you know how to find the slope?

OpenStudy (anonymous):

yes

OpenStudy (lalaly):

can you find the slope of this equation?

OpenStudy (anonymous):

yes the first one is 3...

OpenStudy (anonymous):

LOL i mean the slope is 3

OpenStudy (lalaly):

its actually -3 because (8-5)/(2-3) = 3/-1 = -3

OpenStudy (anonymous):

oh okay so -3. But how do I write the equation?

OpenStudy (lalaly):

now use the equation y-h = m(x-k) (h,k) is any one of the points you have here

OpenStudy (lalaly):

sorry (k,h)

OpenStudy (lalaly):

and m is the slope we found

OpenStudy (anonymous):

THANKS(:

OpenStudy (lalaly):

your welcome:)